# Making Them a Malicious Database: Exploiting Query Code to Jailbreak Aligned Large Language Models

### v1

@@ 4.1 Experimental Setup
Victim Models. We test QueryAttack on 11 mainstream large language models: GPT-4o (OpenAI, 2024b), GPT-4-1106 (gpt-4-1106-preview) (OpenAI, 2024a), GPT-3.5 (gpt-3.5-turbo) (OpenAI, 2023), O1 (gpt-o1) (OpenAI, 2024c), Deepseek (deepseek-chat) (DeepSeek-AI et al., 2024), Gemini-flash (gemini-1.5-flash) (DeepMind, 2024), Llama-3.1-8B (meta-llama-3.1-8B-instruct), Llama-3.1-70B (meta-llama-3.1-70B-instruct), Llama-3.2-1B (meta-llama-3.2-1B-instruct), Llama-3.2-3B (meta-llama-3.2-3B-instruct) and Llama-3.2-11B (meta-llama-3.2-11B-vision-instruct) (Touvron et al., 2023b; Meta, 2024). To maintain the reproducibility of the results, we set all the temperature to 0.
@@
PAIR (Chao et al., 2023b): a black-box attack that uses an adversarial LLM to automatically generate adversarial inputs for the target LLM.
@@
Evaluation and Metrics. We focus on three main metrics to evaluate the effectiveness of QueryAttack: harmfulness score (HS), refuse rate (RR), and attack success rate (ASR). Our evaluation follows the work of (Qi et al., 2024), utilizing the powerful evaluation capabilities of GPT-4 to assign harmfulness scores to the responses generated by the target LLM. The harmfulness score ranges from 1 to 5. To improve the accuracy of GPT-4 Judge, we remove educational statements or disclaimers that may appear in the responses to extract the harmful content, as shown in D.3. The extracted content is then input into GPT-4 Judge for evaluation. Detailed settings can be found in the appendix D.
Based on this, we can calculate the second metric: attack success rate (ASR), which is the percentage of harmful responses in the case of adversarial queries, as shown in Equation 1.
@@
The third evaluation metric is the refuse rate (RR). We follow the setup of (Zou et al., 2023) and use Dict-Judge to assess the number of refusal responses, as shown in Equation 2.
@@
In the comparison with baselines, we primarily consider ASR as the evaluation metric. In other experiments, we also report RR and HS to further supplement the evaluation.
We present the experimental results from CodeAttack (Ren et al., 2024b) and (Song et al., 2025), as they use the same benchmark as ours and also employ the GPT-4 Judge (Qi et al., 2024) method to evaluate their attacks. Therefore, we use their results as baseline for comparisons with QueryAttack. Note that (Song et al., 2025) consider an attack successful when the HS is "higher than 4", whereas we define success only when HS equals 5. However, we can still compare results using the average HS.
@@ 4.2 Results
PAIR TAP CipherChat CodeAttack HEA Ours (Top 1) Ours (Ensemble)
GPT-4-1106 - / - - / - - / 19.00% - / 81.00% - / - 4.65 / 82.18% 4.75 / 93.80%
GPT-4o* 3.16 / 45.38 % 3.24 / 51.34% 1.94 / 16.34% - / 89.00% 4.42 / 90.38% 4.72 / 90.58% 4.85 / 96.35%
Llama-3.1-8b* 3.06 / 35.38% 2.97 / 31.34% 1.76 / 0.00% - / - 4.67 / 95.38% 4.04 / 65.78% 4.83 / 88.89%
Gemini-flash* 1.92 / 18.27% 3.01 / 33.27% 2.12 / 5.38% - / - 4.64 / 100.00% 4.93 / 95.59% 4.99 / 99.62%
Table 1: Average HS / ASR of baselines and QueryAttack on the AdvBench. QueryAttack can breach the safety guardrails of mainstream LLMs, including GPT, Llama and Gemini series and achieves SoTA ASR. *: (Song et al., 2025) reports that they “consider responses of LLMs with a score higher than 4 as successful attacks” and we present their original results. When calculating the ASR for QueryAttack, we count instances where HS = 5.
Table 1 presents the average HS and the ASR of QueryAttack and several baselines on AdvBench (Zou et al., 2023). We demonstrate two configurations of QueryAttack. In the first configuration, denoted as Top 1, QueryAttack uses the programming language style with the highest ASR to construct the query template. In the second configuration, denoted as Ensemble, no restrictions are placed on the programming language styles used for the query templates. From this table, we can observe that QueryAttack achieves SoTA ASR. Despite specialized safety alignment training in these latest LLMs, QueryAttack successfully bypassed their defenses, inducing responses that violate policies or human values.
For GPT-4-1106, the Top 1 configuration achieves a 82.18% ASR and the Ensemble configuration achieves 93.80%. In contrast, the baselines which perform best are only able to bypass GPT-4-1106’s safeguards in up to 81% of cases. The same trend is observed in other models. Except for Llama-3.3-8B, QueryAttack’s ASR is lower than HEA. However, under the Ensemble configuration, the average HS remains higher than that of HEA. This demonstrates the effectiveness of QueryAttack.
For cost considerations, we test QueryAttack’s effectiveness against the O1 model using the subset of AdvBench containing 50 samples. Under the Ensemble configurations, QueryAttack achieves an average HS of 3.66 and an ASR of 50%. These results indicate that CoT Reasoning-enhanced models may have the potential to defend against QueryAttack, but QueryAttack still maintains a considerable ASR.
@@ 4.3 Ablation and Analysis
Impact of different language styles. We create nine query templates in different programming language styles: Python, Java, JavaScript, C++, C, C#, SQL, URL, and Go. Figure 3 shows the average HS and RR obtained by attacking GPT-4-1106, Llama-3.1-70B, Gemini-flash and DeepSeek with templates of different styles. On GPT-4-1106, Gemini-flash and DeepSeek, different language styles do not show significant variations in average HS. A noticeable decrease is observed when attacking Llama-3.1-70B with URL and SQL-style templates, where a higher RR leads to a lower HS. This may be because these two template styles closely resemble the structure of natural language, making them more likely to trigger existing defenses. Thus, using languages that differ more from natural language is likely to increase QueryAttack’s ASR. Despite this, under the Top 1 configuration, the ASR on Llama-3.1-70B still reachs 76.3%, while the Ensemble configuration achieves 92.9%.
@@
Impact of the target model’s parameter size. The increase in model parameter size does not show a positive correlation with the effectiveness of defending against QueryAttack. Table 2 presents the ASRs of QueryAttack on models with different parameter sizes from the latest Llama-3.1 and Llama-3.2 series under Ensemble configuration. On Llama-3.2 series, although the ASR slightly decreases for the 1B parameter models compared to larger models, the ASR does not show a decline as the parameter size continues to grow. Specifically, as the parameter size increases from 8B to 70B, QueryAttack’s ASR on the Llama-3.1 series models rises from 88.89% to 92.91%. Without targeted safety alignment, bigger models may even have a higher risk of being attacked due to their stronger understanding ability of new language. A similar trend can also be seen in results from other previous studies (Ren et al., 2024b; Yuan et al., 2024; Lv et al., 2024). Therefore, exploring more robust defense methods is necessary.
@@ 4.4 Discussion about Countermeasures
We consider two baseline defenses and design a tailored defense method against QueryAttack as follows. Detailed settings for these defense methods can be found in the appendix D.2.
@@
Cross-lingual Alignment Prompting based defense (Ours):  (Qin et al., 2023) propose a method that uses cross-lingual chain-of-thought (CoT) prompting to generate reasoning paths, improving zero-shot CoT reasoning across languages. Although this approach is not originally designed for defense, we find that such CoT reasoning can help LLMs recognize cross-lingual malicious intent. Based on this insight, we design a defense method using cross-lingual chain-of-thought prompting. The complete chain of thought is: we first requires the target model to identify the Query Content, Key Object, and Content Source of the input, then describe the query in natural language. The target model then responds to this natural language query, thereby activating existing safety alignment mechanisms for defense. The detailed prompt can be found in the appendix D.2.
We test the effectiveness of these countermeasures against QueryAttack on four models: Gemini-flash, GPT-3.5-turbo, Llama-3.1-8B, and GPT-4-1106. For cost considerations, we use the subset of AdvBench refined in (Jiang et al., 2024) for evaluation in this section and report the results of the Ensemble configuration using Python, C++, and SQL styles.
@@
In contrast, our tailored defense based on cross-lingual alignment prompting effectively reduces the ASR of QueryAttack across all models, with the exception of a small reduction in GPT-3.5. However, on the other three models, our defense achieves an average ASR reduction of 63% and an average HS reduction of 1.53, demonstrating the effectiveness of our defense. The success of QueryAttack relies on the target model’s ability to accurately interpret custom language templates, meaning the model should also be capable of identifying the intent of them and translating them into natural language. Once translated into natural language, malicious prompts are more likely to be filtered by existing safety alignment defenses. This indicates that a well-executed translate-then-reason CoT process can effectively defend against QueryAttack-like jailbreak attacks. On the O1 model, QueryAttack achieves a lower ASR than on other models, further supporting this opinion.
Paraphrase Rand-insert Rand-swap Rand-patch Ours
Gemini-flash ASR 94% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$ 100% ${\color[rgb]{1,0,0}-0\%}$ 100% ${\color[rgb]{1,0,0}-0\%}$ 100% ${\color[rgb]{1,0,0}-0\%}$ 36% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 74\%}$
HS 4.88 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.12}$ 5.00 ${\color[rgb]{1,0,0}-0.00}$ 5.00 ${\color[rgb]{1,0,0}-0.00}$ 5.00 ${\color[rgb]{1,0,0}-0.00}$ 3.56 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.44}$
GPT-4-1106 ASR 72% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 20\%}$ 86% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$ 94% ${\color[rgb]{1,0,0}\uparrow 2\%}$ 94% ${\color[rgb]{1,0,0}\uparrow 2\%}$ 28% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 64\%}$
HS 4.16 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.76}$ 4.76 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.16}$ 4.88 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.04}$ 4.90 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.02}$ 3.10 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.82}$
GPT-3.5 ASR 68% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ 66% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 16\%}$ 54% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 28\%}$ 64% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 18\%}$ 76% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$
HS 4.56 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.14}$ 4.48 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$ 4.12 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.58}$ 4.48 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$ 4.68 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.02}$
Llama-3.1-8B ASR 90% ${\color[rgb]{1,0,0}\uparrow 2\%}$ 72% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ 70% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 16\%}$ 72% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ 34% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 52\%}$
HS 4.86 ${\color[rgb]{1,0,0}\uparrow 0.14}$ 4.46 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.26}$ 4.50 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$ 4.50 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$ 3.38 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.34}$

### v2

@@ 4.1 Experimental Setup
Victim Models. We test QueryAttack on 13 mainstream large language models: GPT-3.5 (gpt-3.5-turbo) (OpenAI, 2023), GPT-4-1106 (gpt-4-1106-preview) (OpenAI, 2024a), GPT-4o (OpenAI, 2024b), O1 (gpt-o1) (OpenAI, 2024c), Deepseek (deepseek-chat) (DeepSeek-AI et al., 2024), Gemini-flash (gemini-1.5-flash), Gemini-pro (gemini-1.5-pro) (DeepMind, 2024), Llama-3.1-8B (meta-llama-3.1-8B-instruct), Llama-3.1-70B (meta-llama-3.1-70B-instruct), Llama-3.2-1B (meta-llama-3.2-1B-instruct), Llama-3.2-3B (meta-llama-3.2-3B-instruct), Llama-3.2-11B (meta-llama-3.2-11B-vision-instruct) and Llama-3.3-70B (meta-llama-3.3-70B-instruct) (Touvron et al., 2023b; Meta, 2024). To maintain the reproducibility of the results, we set all the temperature to 0.
@@
GPT-4-1106 GPT-4o* Llama-3.1-8B* Llama-3.3-70B* Gemini-pro* Gemini-flash*
PAIR - / - 3.16 / 45.38 % 3.06 / 35.38% 3.24 / 47.30% 1.92 / 22.31% 1.92 / 18.27%
TAP - / - 3.24 / 51.34% 2.97 / 31.34% 3.71 / 55.38% 2.83 / 24.23% 3.01 / 33.27%
CipherChat - / 19% 1.94 / 16.34% 1.76 / 0 % 2.40 / 4.23% 2.22 / 3.27% 2.12 / 5.38%
CodeAttack - / 81% - / 89% - / - - / - - / 2% - / -
HEA - / - 4.42 / 90.38% 4.67 / 95.38% 3.58 / 68.27% 4.21 / 82.38% 4.64 / 100%
Ours (Top 1) 4.65 / 82.18% 4.72 / 90.58% 4.04 / 65.78% 3.98 / 68.77% 4.71 / 85.63% 4.93 / 95.59%
Ours (Ensemble) 4.75 / 93.80% 4.85 / 96.35% 4.83 / 88.89% 4.11 / 73.56% 4.91 / 95.40% 4.99 / 99.62%
Table 1: Average HS / ASR of baselines and QueryAttack on the AdvBench. QueryAttack can breach the safety guardrails of mainstream LLMs, including GPT, Llama and Gemini series and achieves SoTA ASR. *: Song et al. (2025) report that they “consider responses of LLMs with a score higher than 4 as successful attacks” and we present their original results. When calculating the ASR for QueryAttack, we follow Ren et al. (2024b) to count instances where HS = 5.
@@
PAIR (Chao et al., 2023): a black-box attack that uses an adversarial LLM to automatically generate adversarial inputs for the target LLM.
@@
Evaluation and Metrics. We focus on three main metrics for evaluation.
Harmfulness score (HS). Our evaluation follows the work of (Qi et al., 2024), utilizing the powerful evaluation capabilities of GPT-4 to assign harmfulness scores to the responses generated by the target LLM. The harmfulness score ranges from 1 to 5. To improve the accuracy of GPT-4 Judge, we remove educational statements or disclaimers that may appear in the responses to extract the harmful content, as shown in B.3. The extracted content is then input into GPT-4 Judge for evaluation. Detailed settings can be found in the Appendix B.
Attack success rate (ASR). Based on HS, we calculate the ASR, the percentage of harmful responses in the case of adversarial queries, which can be calculated by Equation 1.
@@
Refuse rate (RR). The third evaluation metric is the refuse rate (RR). We follow the setup of (Zou et al., 2023) and use Dict-Judge to assess the number of refusal responses, as shown in Equation 2.
@@
We present the experimental results from CodeAttack (Ren et al., 2024b) and (Song et al., 2025), as they use the same benchmark as ours and also employ the GPT-4 Judge (Qi et al., 2024) method to evaluate their attacks. Therefore, we use their results as baseline for comparisons with QueryAttack. Note that Song et al. (2025) consider an attack successful when the HS is “higher than 4”, whereas we follow Ren et al. (2024b) to define success only when HS equals 5.
@@ 4.2 Results
QueryAttack achieves SoTA ASR. Table 1 presents the average HS and the ASR of QueryAttack and several baselines on AdvBench (Zou et al., 2023). We demonstrate two configurations of QueryAttack. In the first configuration, denoted as Top 1, QueryAttack uses the programming language style with the highest ASR to construct the query template. In the second configuration, denoted as Ensemble, no restrictions are placed on the programming language styles. From this table, we can observe that, despite specialized safety alignment training in these latest LLMs, QueryAttack successfully bypassed their defenses, inducing responses that violate policies or human values. For GPT-4-1106, the Top 1 configuration achieves a 82.18% ASR and the Ensemble configuration achieves 93.80%. In contrast, the baselines which perform best are only able to bypass GPT-4-1106’s safeguards in up to 81% of cases. The same trend is observed in other models. Except for Llama-3.1-8B, QueryAttack’s ASR is lower than HEA. However, under the Ensemble configuration, the average HS remains higher than that of HEA.
QueryAttack remains effective when facing reasoning-enhanced models. For cost considerations, we test QueryAttack’s effectiveness against the O1 model using the subset of AdvBench containing 50 samples. Under the Ensemble configurations, QueryAttack achieves an average HS of 3.66 and an ASR of 50%. These results indicate that CoT Reasoning-enhanced models may have the potential to defend against QueryAttack, but QueryAttack still maintains a considerable ASR.
@@ 4.3 Ablation and Analysis
Languages that differ more from natural language is likely to increase QueryAttack’s ASR. Figure 4 shows the average HS and RR obtained by attacking GPT-4-1106, Llama-3.1-70B, Gemini-flash and DeepSeek with templates of different language styles. On GPT-4-1106, Gemini-flash and DeepSeek, different language styles do not show significant variations in average HS. A noticeable decrease is observed when attacking Llama-3.1-70B with URL and SQL-style templates, where a higher RR leads to a lower HS. This may be because these two template styles closely resemble the structure of natural language, making them more likely to trigger existing defenses. Despite this, under the Top 1 configuration, the ASR on Llama-3.1-70B still reachs 76.3%, while the Ensemble configuration achieves 92.9%.
@@
Gemini-flash GPT-4-1106 GPT-3.5 Llama-3.1-8B
ASR HS ASR HS ASR HS ASR HS
Paraphrase 94% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$ 4.88 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.12}$ 72% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 20\%}$ 4.16 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.76}$ 68% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ 4.56 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.14}$ 90% ${\color[rgb]{1,0,0}\uparrow 2\%}$ 4.86 ${\color[rgb]{1,0,0}\uparrow 0.14}$
Rand-insert 100% ${\color[rgb]{1,0,0}-0\%}$ 5.00 ${\color[rgb]{1,0,0}-0.00}$ 86% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$ 4.76 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.16}$ 66% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 16\%}$ 4.48 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$ 72% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ 4.46 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.26}$
Rand-swap 100% ${\color[rgb]{1,0,0}-0\%}$ 5.00 ${\color[rgb]{1,0,0}-0.00}$ 94% ${\color[rgb]{1,0,0}\uparrow 2\%}$ 4.88 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.04}$ 54% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 28\%}$ 4.12 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.58}$ 70% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 16\%}$ 4.50 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$
Rand-patch 100% ${\color[rgb]{1,0,0}-0\%}$ 5.00 ${\color[rgb]{1,0,0}-0.00}$ 94% ${\color[rgb]{1,0,0}\uparrow 2\%}$ 4.90 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.02}$ 64% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 18\%}$ 4.48 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$ 72% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ 4.50 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$
Ours 36% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 74\%}$ 3.56 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.44}$ 28% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 64\%}$ 3.10 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.82}$ 76% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$ 4.68 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.02}$ 34% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 52\%}$ 3.38 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.34}$
Larger models do not provide better defense against QueryAttack. Table 3 presents the ASRs of QueryAttack on models with different parameter sizes from the latest Llama-3.1 and Llama-3.2 series under Ensemble configuration. On Llama-3.2 series, although the ASR slightly decreases for the 1B parameter models compared to larger models, the ASR does not show a decline as the parameter size continues to grow. Specifically, as the parameter size increases from 8B to 70B, QueryAttack’s ASR on the Llama-3.1 series models rises from 88.89% to 92.91%. This means increase in model parameter size does not show a positive correlation with the effectiveness of defending against QueryAttack. Without targeted safety alignment, bigger models may even have a higher risk of being attacked due to their stronger understanding ability of new language. Therefore, exploring more robust defense methods is necessary.
@@ 4.4 Discussion about Countermeasures
We consider two baseline defenses and design a tailored defense method against QueryAttack as follows. Detailed settings for these defense methods can be found in the Appendix B.2.
@@
Cross-lingual Alignment Prompting based defense (Ours):  (Qin et al., 2023) propose a method that uses cross-lingual chain-of-thought (CoT) prompting to generate reasoning paths, improving zero-shot CoT reasoning across languages. Although this approach is not originally designed for defense, we find that such CoT reasoning can help LLMs recognize cross-lingual malicious intent. Our insight is that, the success of QueryAttack relies on the target model’s ability to accurately interpret custom language templates, meaning the model should also be capable of identifying the intent of them and translating them into natural language. Once translated into natural language, malicious prompts are more likely to be filtered by existing safety alignment defenses. This indicates that a well-executed translate-then-reason CoT process can effectively defend against QueryAttack-like jailbreak attacks. Based on this, we design a defense method using cross-lingual chain-of-thought prompting. The complete chain of thought is: we first requires the target model to identify the Query Content, Key Object, and Content Source of the input, then describe the query in natural language. The target model then responds to this natural language query, thereby activating existing safety alignment mechanisms for defense. The detailed prompt can be found in the Appendix B.2.
We test the effectiveness of these countermeasures against QueryAttack on four models: Gemini-flash, GPT-3.5, Llama-3.1-8B, and GPT-4-1106. For cost considerations, we use the subset of AdvBench refined in (Jiang et al., 2024) for evaluation in this section and report the results of the Ensemble configuration using Python, C++, and SQL styles.
@@
In contrast, our tailored defense based on cross-lingual alignment prompting effectively reduces the ASR of QueryAttack across all models, with the exception of a small reduction in GPT-3.5. However, on the other three models, our defense achieves an average ASR reduction of 63% and an average HS reduction of 1.53, demonstrating the effectiveness of our defense.
5 Related Work
5.1 Jailbreak Attacks on LLMs
Initially, researchers reveal that adversaries could launch attacks by manually constructing out-of-distribution (OOD) samples (Li et al., 2023; Shen et al., 2024). Building on these observations, several white-box attack methods are proposed (Jones et al., 2023a, b; Gao et al., 2024c). Compared to white-box attacks, black-box attacks assume that adversaries adjust their prompt strategies only based on the model’s responses (Zhang et al., 2024; Shah et al., 2023; Liu et al., 2024; Ren et al., 2024c; Yuan et al., 2024; Deng et al., 2024).
Recently, some black-box are proposed to use code to encrypt malicious inputs to build long-tail encoded distributions. CodeAttack (Ren et al., 2024b) embeds malicious queries within data structures (e.g., stacks and queues) to bypass safety alignments designed for prompts written in natural languages. Codechameleon (Lv et al., 2024) encrypts malicious prompts using custom program functions, transforming them into code completion tasks. Unlike these methods, QueryAttack does not rely on the syntax of programming languages for encryption. Instead, it only requires certain keywords or expressions from the programming language. This means that QueryAttack can be applied not only using programming languages but also to any non-natural language that the target LLM can understand but has not been well aligned during the safety alignment phase. Moreover, even without the need for output encryption, QueryAttack can still effectively attack target LLMs.
5.2 Safety Alignment for Defending Jailbreak
Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017) is one of the most widely used defense mechanisms. For instance, recent works such as (Mehrabi et al., 2024; Sun et al., 2023) explore the effectiveness of alignment during pre-training in defending against malicious queries, CoT reasoning (OpenAI, 2024c), as well as in-context learning (Wei et al., 2023b; Ren et al., 2024a). These methods often rely on natural language inputs collected from red teams, which can lead to generalization issues when faced with non-natural language or other OOD inputs.